# Advances in Feed-Forward 3D Reconstruction and View Synthesis: A Survey

### v3

@@ Abstract
3D reconstruction and view synthesis are foundational problems in computer vision, graphics, and immersive technologies such as augmented reality (AR), virtual reality (VR), and digital twins. Traditional methods rely on computationally intensive iterative optimization in a complex chain, limiting their applicability in real-world scenarios. Recent advances in feed-forward approaches, driven by deep learning, have revolutionized this field by enabling fast and generalizable 3D reconstruction and view synthesis. This survey offers a comprehensive review of feed-forward techniques for 3D reconstruction and view synthesis, with a taxonomy according to the underlying representation architectures including point cloud, 3D Gaussian Splatting (3DGS), Neural Radiance Fields (NeRF), etc. We examine key tasks such as pose-free reconstruction, dynamic 3D reconstruction, and 3D-aware image and video synthesis, highlighting their applications in digital humans, SLAM, robotics, and beyond. In addition, we review commonly used datasets with detailed statistics, along with evaluation protocols for various downstream tasks. We conclude by discussing open research challenges and promising directions for future work, emphasizing the potential of feed-forward approaches to advance the state of the art in 3D vision. A project page associated with this survey is available at Feed-Forward-3D.
Index Terms:
Feed-forward Model, 3D Reconstruction, Neural Rendering, Radiance Fields, NeRF, 3DGS.
@@
Figure 2: The above figure includes various applications of feed-forward models in 3D reconstruction and Novel View Synthesis (NVS). The samples are adapted from [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
1 Introduction
3D reconstruction and rendering are long-standing and central challenges in computer vision and computer graphics. They enable a wide range of applications, from digital content creation, augmented reality, and virtual reality to robotics, autonomous systems, and digital twins. Traditionally, high-quality 3D reconstruction and view synthesis have relied on optimization-based pipelines such as Structure-from-Motion (SfM) [13] and 3D Gaussian Splatting [14]. However, these methods typically require per-scene optimization, converge slowly, and rely on dense multiview inputs, limiting their practicality in real-world scenarios. In light of this, feed-forward methods have emerged as an important research line in 3D vision.
Feed-forward models are learning-based methods that infer outputs, such as 3D geometry or novel views, in a single forward pass of a neural network, without requiring per-scene iterative optimization in classical methods. Early examples include cost-volume-based Multi-View Stereo [15] and layered representations such as multiplane images (MPI) [16], which demonstrate the potential of learning-based inference over per-scene optimization. In recent years, fueled by breakthroughs in deep learning and neural representations, feed-forward methods [17, 18, 19] have emerged as a transformative alternative in 3D reconstruction and view synthesis. By learning from large scale data, these models enable orders of magnitude faster inference and better generalization compared to classical optimization methods. Their ability to produce predictions in real time makes them particularly attractive for time-sensitive and scalable applications, such as robotic perception and interactive 3D asset creation.
This survey focuses on feed-forward methods developed primarily after the emergence of neural radiance fields (NeRF) [20] in 2020, which catalyzed a rapid evolution in feed-forward approaches, as shown in Fig. 1. We present a comprehensive review of feed-forward methods for 3D reconstruction and view synthesis, with an emphasis on the scene representations, core architectures, and downstream applications that define this fast-evolving area. Specifically, we systematically categorize existing approaches based on their underlying scene representations, which determine how 3D structure and appearance are modeled and rendered. In this way, we identify five major categories: 1) models built on Neural Radiance Fields (NeRF) [20], which leverage volumetric rendering through learned radiance fields; 2) pointmap-based approaches [1], which operate on pixel-aligned 3D pointmaps; 3) 3D Gaussian Splatting (3DGS)-based models [14], which use rasterizable Gaussian primitives for fast and efficient rendering; 4) methods based on other 3D representations like mesh, occupancy and signed distance function (SDF); and 5) 3D-free models, which leverage deep neural networks to synthesize views directly without a explicit 3D representation. For each category, we provide an in-depth analysis of representative and state-of-the-art methods, highlighting their core architectural designs and feature representations.
We also highlight several high-impact 3D vision applications enabled by feed-forward methods, as illustrated in Fig. 2. These methods offer scalable, fast, and generalizable solutions across domains, supporting tasks such as multi-view 3D reconstruction, dynamic 3D Gaussian Splatting, and diffusion-based view synthesis. Beyond reconstruction and view synthesis, they also advance image matching, 3D-aware segmentation, and optical flow estimation. In the broader context of robotics and SLAM, feed-forward models enable efficient real-time scene understanding and tracking, while in digital humans, they support generalizable avatar reconstruction from sparse inputs.
@@ 2 Methods
Figure 3: Representative frameworks of feed-forward NeRF, including methods based on (a) 1D features, (b) 2D features, and (c) & (d) 3D features. The samples are adapted from [21] [22] [23] [10].
@@ 2.1 NeRF
Neural radiance fields (NeRF) [20] have recently gained significant attention for high-quality novel view synthesis using volumetric scene representations and differentiable volume rendering. By leveraging MLPs, NeRF reconstructs 3D scenes from multiview 2D images, enabling the generation of novel views with excellent multiview consistency. However, a major limitation of NeRF is its requirement for per-scene optimization, so it cannot generalize to unseen scenes. To address this, feed-forward approaches have been proposed, where neural networks learn to infer NeRF representations directly from sparse input views, thereby eliminating the need for scene-specific optimization. As a pioneering feed-forward NeRF work, PixelNeRF [19] introduces a conditional NeRF framework that leverages pixel-aligned image features extracted from input images, allowing the model to generalize across diverse scenes and perform novel view synthesis from sparse observations. A large number of follow-ups adopt various techniques for feed-forward NeRF, and we broadly categorize them into the following categories based on feature representations.
@@
Several methods have been proposed to encode a global 1D latent code for NeRF prediction, where the same latent code is shared between all 3D points in a scene. For example, CodeNeRF [21], as illustrated in Fig. 3(a), introduces a disentanglement strategy that jointly learns separate embeddings for texture and shape, along with an MLP conditioned on these embeddings to predict the color and volumetric density of each 3D point. ShaRF [24] introduces latent codes for shape and appearance, which serve as conditioning inputs for NeRF reconstruction. In addition, Shap-E [25] encodes point clouds and RGBA input images into a series of latent vectors that are subsequently used for generating NeRF’s MLPs.
@@
2D feature-based methods typically leverage an image encoder to extract image features of source views and obtain features of arbitrary 3D points by ray projection without relying on 3D intermediate features. For example, GRF [22], as illustrated in Fig. 3(b) projects each 3D point along a camera ray onto source views to extract corresponding multiview features. These features are then aggregated and passed through an MLP to predict color and density. IBRNet [26] follows a similar approach, projecting 3D points onto nearby source views to extract image features that are aggregated across views for NeRF inference. NeRFormer [27] also employs ray-projected features and performs multiview feature aggregation to guide the NeRF prediction. In addition, SRF [28] projects 3D points onto multiview images to construct a stereo feature matrix, which is processed by a 2D CNN to produce view-aligned features for radiance prediction. To provide additional geometric cues for NeRF prediction, GNT [29] introduces a view transformer that uses epipolar constraint to aggregate projected features from multiple views in a geometrically consistent manner. Its successor, GNT-MOVE [30], bridges the view transformer with the Mixture-of-Experts concept from large language models, enhancing its cross-scene generalization capability. Instead of relying on epipolar lines, MatchNeRF [31] explicitly models the correspondence information by computing the similarity between ray-projected features from pairs of nearby source views, using this information as a conditioning input for NeRF prediction. In addition, ContraNeRF [32] introduces geometry-aware contrastive learning [33] to learn multiview consistent 2D features to mitigate synthetic-to-real generalization issues.
@@
3D Volume Features. Inspired by multiview stereo (MVS) [15, 34, 35], MVSNeRF [23] constructs 3D cost volumes that store the matching costs of pixels/features across multiple images from input images (usually along depth values). These cost volumes are used to generate a neural scene encoding volume, which stores per-voxel features representing both local geometry and appearance, as shown in Fig. 3(c). For any 3D point, its features are obtained via trilinear interpolation from the encoding volume and then decoded by an MLP to predict the corresponding density and color. To improve rendering quality in both fine-detail areas and occluded regions, GeoNeRF [36] extends MVSNeRF by first constructing cascaded 3D cost volumes for each source view, followed by an attention-based volume aggregation across views. Similarly addressing occlusions, NeuRay [37] leverages cost volumes to predict the visibility of 3D points, which can identify feature inconsistencies caused by occlusion. WaveNeRF [38] further enhances geometry reconstruction by integrating wavelet frequency volumes into the MVS pipeline to preserve high-frequency details.
@@
For efficient rendering, ENeRF [39] samples a limited number of points near the scene surface, guided by coarse scene geometry predicted from a cascaded cost volume, thereby speeding up rendering; Instead of constructing cost volumes for all reference input views, MuRF [40] efficiently constructs a target view frustum volume to aggregate information from the input images; To improve the quality of geometry estimation, GeFu [41] introduces an adaptive cost aggregation module that learns to reweight contributions from different source views, enabling more accurate cost volume construction.
@@
3D Triplane Features. The 3D triplane serves as an efficient volumetric representation [42, 43], making it highly compatible with feed-forward models. Specifically, Large Reconstruction Model (LRM) [10] employs a large transformer-based encoder-decoder architecture and directly regresses a feature triplane representation as shown in Fig. 3(d), enabling NeRF prediction from triplane features. Pf-LRM [44] extends LRM to a pose-free setting, which jointly reconstructs the triplane NeRF representations and predicts relative camera poses. TripoSR [45] further advances LRM by carefully curating a subset of Objaverse [46] for training, along with improved architectures and training strategies. To address the scarcity, licensing constraints, and inherent biases of 3D data, LRM-Zero [47] introduces the Zeroverse dataset and performs training entirely on synthesized data. Beyond reconstruction alone, several methods integrate LRM with diffusion models. For example, Instant3D [48] first leverages a fine-tuned 2D diffusion model [49] to generate 4-view images from a text prompt and then uses a transformer-based large reconstruction model to predict a NeRF; DMV3D [50] incorporates LRM into multiview diffusion, which gradually reconstructs a clean triplane NeRF representation from noisy multiview images in the diffusion process.
@@
In addition to the methods discussed above, several approaches have explored feed-forward NeRF reconstruction with other types of features. For example, VisionNeRF [51] proposes to leverage vision transformer [52] and convolutional networks to extract global 1D features and 2D image features, respectively, and constructs a multi-level feature map that serves as the conditioning inputs of NeRF prediction to enhance rendering quality, particularly in occluded regions. MINE [53] integrates NeRF and multiplane image (MPI) [54] representations to enable generalizable, occlusion-aware 3D reconstruction from a single image.
@@ 2.2 Pointmap
Figure 4: The framework of feed-forward pointmap reconstruction. It also supports broader tasks, such as camera estimation, depth estimation, and point tracking, within the same framework. The samples are adapted from [3].
@@
Pointmaps [1, 55, 56, 57, 58], encode scene geometry, pixel-to-scene correspondences, and viewpoint relationships, allowing for camera poses, depths, and explicit 3D primitive estimation as shown in Fig. 4. The pioneering feed-forward pointmap reconstruction method DUSt3R [1] learns a transformer-based encoder-decoder to directly output two pixel-aligned pointmaps from image pairs without posed cameras, enabling dense unconstrained stereo 3D reconstruction. The follow-up work, MASt3R [4], improves DUSt3R by introducing local feature matching.
@@
To handle more views, Fast3R [59] builds on DUSt3R and designs a global fusion transformer to process multiview inputs simultaneously. MV-DUSt3R [60] instead leverages multiview decoder blocks to learn both the reference-to-source and source-to-source view relationships, thereby extending DUSt3R to a multiview setting. SLAM3R [61] introduces an Image-to-Points module that enables simultaneous processing of multiview inputs, effectively enhancing reconstruction quality without sequential reconstruction.
@@
Notably, several works incorporate a memory mechanism that incrementally processes inputs and updates a scene’s latent state, progressively adding points to a canonical 3D representation. Spann3R [62] introduces a spatial memory network that enables multiview inputs while improving efficiency, removing the need for global alignment. Similarly, MUSt3R [63] extends the DUSt3R architecture with a symmetric design and a memory mechanism, effectively reducing computational complexity when processing multiview inputs. CUT3R [3] proposes a Continuous Updating Transformer that simultaneously updates the state with new information and retrieves the information stored in the state. This formulation is general, supporting video and photo collections as well as static and dynamic scenes. However, with the increased number of processed frames, memory-based methods face capacity constraints, which can result in the degradation or loss of information from earlier frames. To address this issue, Point3R [64] takes inspiration from the human memory mechanism and proposes a spatial pointer memory, where each pointer is anchored at a 3D position and links to a dynamically evolving spatial feature. Driv3R [65] extends the memory mechanism to support efficient temporal integration, enabling large-scale dynamic scene reconstruction from multiview input sequences.
@@
yuelei here In addition, several methods are proposed to develop new SfM pipelines for efficient 3D reconstruction. Specifically, Light3R-SfM [66] replaces optimization-based global alignment with a learnable latent alignment module, enabling the efficient SfM and 3D reconstruction. Regist3R [67] introduces a stereo foundation model to build a scalable incremental SfM pipeline for efficient 3D reconstruction.
@@
To facilitate accurate 3D reconstruction, Pow3R [68] flexibly integrates available priors at test time, such as camera intrinsics, sparse or dense depth, or relative poses, as lightweight and diverse conditioning. In contrast, Rig3R [69] exploits the rig metadata as conditions to improve both the camera pose estimation and 3D reconstruction. Meanwhile, MoGe [70] improves geometry learning by replacing DUSt3R’s scale-invariant pointmaps with affine-invariant ones, and further introduces a global alignment solver to correct scale and shift errors in affine-invariant pointmaps. Similarly focused on geometric consistency, Test3R [71] takes advantage of test time training to improve the geometric consistency of pointmaps. AerialMegaDepth [72] instead focuses on aerial-ground geometric reconstruction from a data perspective.
@@
As a promising and powerful foundation for 3D reconstruction, VGGT [2] presents a large feed-forward transformer-based architecture that directly predicts all essential 3D attributes, such as camera intrinsics and extrinsics, point maps, depth maps, and 3D point tracks, without the need for post-processing, leading to state-of-the-art 3D point and camera pose reconstruction. VGGT-Long [73] extends VGGT to handle kilometer-scale sequences through a chunk-and-align pipeline, effectively mitigating its memory constraints in long-sequence handling.
@@
3D Gaussian Splatting (GS) [14] is a recent advance for efficient 3D reconstruction and rendering built on rasterization. Despite its high fidelity in reconstruction, 3DGS requires per-scene optimization, which limits its training efficiency and generalization capabilities. Recently, feed-forward 3DGS reconstruction methods have been developed by leveraging neural networks to directly predict Gaussian parameters, eliminating the need for per-scene optimization. We categorize these methods based on the representation of predicted Gaussian outputs: 2D map and 3D volume.
Figure 5: Representative frameworks with different outputs of 3D Gaussian representations, including Gaussian Map (e.g., Image Splatter and pointmap) and Gaussian Volume (e.g., Triplane and 3D Volume). The samples are adapted from [74] [75] [76] [77].

### v4

@@ Abstract
[…]
1 Introduction
3D reconstruction and rendering are long-standing and central challenges in computer vision and computer graphics. They enable a wide range of applications, from digital content creation, augmented reality, and virtual reality to robotics, autonomous systems, and digital twins. Traditionally, high-quality 3D reconstruction and view synthesis have relied on optimization-based pipelines such as Structure-from-Motion (SfM) [SF16] and 3D Gaussian Splatting [KKLD23]. However, these methods typically require per-scene optimization, converge slowly, and rely on dense multiview inputs, limiting their practicality in real-world scenarios. In light of this, feed-forward methods have emerged as an important research line in 3D vision.
Feed-forward models are learning-based methods that infer outputs, such as 3D geometry or novel views, in a single forward pass of a neural network, without requiring per-scene iterative optimization in classical methods. Early examples include cost-volume-based Multi-View Stereo [YLL18] and layered representations such as multiplane images (MPI) [ZTF18], which demonstrate the potential of learning-based inference over per-scene optimization. In recent years, fueled by breakthroughs in deep learning and neural representations, feed-forward methods [FSG17, WZX16, YYTK21] have emerged as a transformative alternative in 3D reconstruction and view synthesis. By learning from large scale data, these models enable orders of magnitude faster inference and better generalization compared to classical optimization methods. Their ability to produce predictions in real time makes them particularly attractive for time-sensitive and scalable applications, such as robotic perception and interactive 3D asset creation.
This survey focuses on feed-forward methods developed primarily after the emergence of neural radiance fields (NeRF) [MST21] in 2020, which catalyzed a rapid evolution in feed-forward approaches, as shown in Fig. 1. We present a comprehensive review of feed-forward methods for 3D reconstruction and view synthesis, with an emphasis on the scene representations, core architectures, and downstream applications that define this fast-evolving area. Specifically, we systematically categorize existing approaches based on their underlying scene representations, which determine how 3D structure and appearance are modeled and rendered. In this way, we identify five major categories: 1) models built on Neural Radiance Fields (NeRF) [MST21], which leverage volumetric rendering through learned radiance fields; 2) pointmap-based approaches [WLC24a], which operate on pixel-aligned 3D pointmaps; 3) 3D Gaussian Splatting (3DGS)-based models [KKLD23], which use rasterizable Gaussian primitives for fast and efficient rendering; 4) methods based on other 3D representations like mesh, occupancy and signed distance function (SDF); and 5) 3D-free models, which leverage deep neural networks to synthesize views directly without a explicit 3D representation. For each category, we provide an in-depth analysis of representative and state-of-the-art methods, highlighting their core architectural designs and feature representations.
We also highlight several high-impact 3D vision applications enabled by feed-forward methods, as illustrated in Fig. Advances in Feed-Forward 3D Reconstruction and View Synthesis: A Survey. These methods offer scalable, fast, and generalizable solutions across domains, supporting tasks such as multi-view 3D reconstruction, dynamic 3D Gaussian Splatting, and diffusion-based view synthesis. Beyond reconstruction and view synthesis, they also advance image matching, 3D-aware segmentation, and optical flow estimation. In the broader context of robotics and SLAM, feed-forward models enable efficient real-time scene understanding and tracking, while in digital humans, they support generalizable avatar reconstruction from sparse inputs.
@@ 2 Methods
Figure 2: Representative frameworks of feed-forward NeRF, including methods based on (a) 1D features, (b) 2D features, and (c) & (d) 3D features. The samples are adapted from [JA21] [TY21] [CXZ21] [HZG23].
@@ 2.1 NeRF
Neural radiance fields (NeRF) [MST21] have recently gained significant attention for high-quality novel view synthesis using volumetric scene representations and differentiable volume rendering. By leveraging MLPs, NeRF reconstructs 3D scenes from multiview 2D images, enabling the generation of novel views with excellent multiview consistency. However, a major limitation of NeRF is its requirement for per-scene optimization, so it cannot generalize to unseen scenes. To address this, feed-forward approaches have been proposed, where neural networks learn to infer NeRF representations directly from sparse input views, thereby eliminating the need for scene-specific optimization. As a pioneering feed-forward NeRF work, PixelNeRF [YYTK21] introduces a conditional NeRF framework that leverages pixel-aligned image features extracted from input images, allowing the model to generalize across diverse scenes and perform novel view synthesis from sparse observations. A large number of follow-ups adopt various techniques for feed-forward NeRF, and we broadly categorize them into the following categories based on feature representations.
@@
Several methods have been proposed to encode a global 1D latent code for NeRF prediction, where the same latent code is shared between all 3D points in a scene. For example, CodeNeRF [JA21], as illustrated in Fig. 2(a), introduces a disentanglement strategy that jointly learns separate embeddings for texture and shape, along with an MLP conditioned on these embeddings to predict the color and volumetric density of each 3D point. ShaRF [RMBF21] introduces latent codes for shape and appearance, which serve as conditioning inputs for NeRF reconstruction. In addition, Shap-E [JN23] encodes point clouds and RGBA input images into a series of latent vectors that are subsequently used for generating NeRF’s MLPs.
@@
2D feature-based methods typically leverage an image encoder to extract image features of source views and obtain features of arbitrary 3D points by ray projection without relying on 3D intermediate features. For example, GRF [TY21], as illustrated in Fig. 2(b) projects each 3D point along a camera ray onto source views to extract corresponding multiview features. These features are then aggregated and passed through an MLP to predict color and density. IBRNet [WWG21] follows a similar approach, projecting 3D points onto nearby source views to extract image features that are aggregated across views for NeRF inference. NeRFormer [RSH21] also employs ray-projected features and performs multiview feature aggregation to guide the NeRF prediction. In addition, SRF [CBLPM21] projects 3D points onto multiview images to construct a stereo feature matrix, which is processed by a 2D CNN to produce view-aligned features for radiance prediction. To provide additional geometric cues for NeRF prediction, GNT [WCC22] introduces a view transformer that uses epipolar constraint to aggregate projected features from multiple views in a geometrically consistent manner. Its successor, GNT-MOVE [CLW23], bridges the view transformer with the Mixture-of-Experts concept from large language models, enhancing its cross-scene generalization capability. Instead of relying on epipolar lines, MatchNeRF [CXW23] explicitly models the correspondence information by computing the similarity between ray-projected features from pairs of nearby source views, using this information as a conditioning input for NeRF prediction. In addition, ContraNeRF [YHL23] introduces geometry-aware contrastive learning [CKNH20] to learn multiview consistent 2D features to mitigate synthetic-to-real generalization issues.
@@
3D Volume Features. Inspired by multiview stereo (MVS) [YLL18, GFZ20, CXZ20], MVSNeRF [CXZ21] constructs 3D cost volumes that store the matching costs of pixels/features across multiple images from input images (usually along depth values). These cost volumes are used to generate a neural scene encoding volume, which stores per-voxel features representing both local geometry and appearance, as shown in Fig. 2(c). For any 3D point, its features are obtained via trilinear interpolation from the encoding volume and then decoded by an MLP to predict the corresponding density and color. To improve rendering quality in both fine-detail areas and occluded regions, GeoNeRF [JLF22] extends MVSNeRF by first constructing cascaded 3D cost volumes for each source view, followed by an attention-based volume aggregation across views. Similarly addressing occlusions, NeuRay [LPL22] leverages cost volumes to predict the visibility of 3D points, which can identify feature inconsistencies caused by occlusion. WaveNeRF [XZZ23] further enhances geometry reconstruction by integrating wavelet frequency volumes into the MVS pipeline to preserve high-frequency details.
@@
For efficient rendering, ENeRF [LPX22] samples a limited number of points near the scene surface, guided by coarse scene geometry predicted from a cascaded cost volume, thereby speeding up rendering; Instead of constructing cost volumes for all reference input views, MuRF [XCC24] efficiently constructs a target view frustum volume to aggregate information from the input images; To improve the quality of geometry estimation, GeFu [LYS24] introduces an adaptive cost aggregation module that learns to reweight contributions from different source views, enabling more accurate cost volume construction.
@@
3D Triplane Features. The 3D triplane serves as an efficient volumetric representation [PNM20, CLC22], making it highly compatible with feed-forward models. Specifically, Large Reconstruction Model (LRM) [HZG23] employs a large transformer-based encoder-decoder architecture and directly regresses a feature triplane representation as shown in Fig. 2(d), enabling NeRF prediction from triplane features. Pf-LRM [WTB23] extends LRM to a pose-free setting, which jointly reconstructs the triplane NeRF representations and predicts relative camera poses. TripoSR [TPL24] further advances LRM by carefully curating a subset of Objaverse [DSS23] for training, along with improved architectures and training strategies. To address the scarcity, licensing constraints, and inherent biases of 3D data, LRM-Zero [XBS24] introduces the Zeroverse dataset and performs training entirely on synthesized data. Beyond reconstruction alone, several methods integrate LRM with diffusion models. For example, Instant3D [LTZ23] first leverages a fine-tuned 2D diffusion model [PEL23] to generate 4-view images from a text prompt and then uses a transformer-based large reconstruction model to predict a NeRF; DMV3D [XTL23] incorporates LRM into multiview diffusion, which gradually reconstructs a clean triplane NeRF representation from noisy multiview images in the diffusion process.
@@
In addition to the methods discussed above, several approaches have explored feed-forward NeRF reconstruction with other types of features. For example, VisionNeRF [LLL23] proposes to leverage vision transformer [DBK20] and convolutional networks to extract global 1D features and 2D image features, respectively, and constructs a multi-level feature map that serves as the conditioning inputs of NeRF prediction to enhance rendering quality, particularly in occluded regions. MINE [LFS21] integrates NeRF and multiplane image (MPI) [TS20] representations to enable generalizable, occlusion-aware 3D reconstruction from a single image.
@@ 2.2 Pointmap
Figure 3: The framework of feed-forward pointmap reconstruction. It also supports broader tasks, such as camera estimation, depth estimation, and point tracking, within the same framework. The samples are adapted from [WZH25].
@@
Pointmaps [WLC24a, BKN17, BR18, BR21, DWZ22], encode scene geometry, pixel-to-scene correspondences, and viewpoint relationships, allowing for camera poses, depths, and explicit 3D primitive estimation as shown in Fig. 3. The pioneering feed-forward pointmap reconstruction method DUSt3R [WLC24a] learns a transformer-based encoder-decoder to directly output two pixel-aligned pointmaps from image pairs without posed cameras, enabling dense unconstrained stereo 3D reconstruction. The follow-up work, MASt3R [LCR24], improves DUSt3R by introducing local feature matching.
@@
To handle more views, Fast3R [YSL25] builds on DUSt3R and designs a global fusion transformer to process multiview inputs simultaneously. MV-DUSt3R [TFW25] instead leverages multiview decoder blocks to learn both the reference-to-source and source-to-source view relationships, thereby extending DUSt3R to a multiview setting. SLAM3R [LDW25] introduces an Image-to-Points module that enables simultaneous processing of multiview inputs, effectively enhancing reconstruction quality without sequential reconstruction.
@@
Notably, several works incorporate a memory mechanism that incrementally processes inputs and updates a scene’s latent state, progressively adding points to a canonical 3D representation. Spann3R [WA24] introduces a spatial memory network that enables multiview inputs while improving efficiency, removing the need for global alignment. Similarly, MUSt3R [CSA25] extends the DUSt3R architecture with a symmetric design and a memory mechanism, effectively reducing computational complexity when processing multiview inputs. CUT3R [WZH25] proposes a Continuous Updating Transformer that simultaneously updates the state with new information and retrieves the information stored in the state. This formulation is general, supporting video and photo collections as well as static and dynamic scenes. However, with the increased number of processed frames, memory-based methods face capacity constraints, which can result in the degradation or loss of information from earlier frames. To address this issue, Point3R [WZZL25] takes inspiration from the human memory mechanism and proposes a spatial pointer memory, where each pointer is anchored at a 3D position and links to a dynamically evolving spatial feature. Driv3R [FZD24] extends the memory mechanism to support efficient temporal integration, enabling large-scale dynamic scene reconstruction from multiview input sequences.
@@
In addition, several methods are proposed to develop new SfM pipelines for efficient 3D reconstruction. Specifically, Light3R-SfM [EZLT25] replaces optimization-based global alignment with a learnable latent alignment module, enabling the efficient SfM and 3D reconstruction. Regist3R [LLQD25] introduces a stereo foundation model to build a scalable incremental SfM pipeline for efficient 3D reconstruction.
@@
To facilitate accurate 3D reconstruction, Pow3R [JWL25] flexibly integrates available priors at test time, such as camera intrinsics, sparse or dense depth, or relative poses, as lightweight and diverse conditioning. In contrast, Rig3R [LKC25] exploits the rig metadata as conditions to improve both the camera pose estimation and 3D reconstruction. Meanwhile, MoGe [WXD25] improves geometry learning by replacing DUSt3R’s scale-invariant pointmaps with affine-invariant ones, and further introduces a global alignment solver to correct scale and shift errors in affine-invariant pointmaps. Similarly focused on geometric consistency, Test3R [YSW25] takes advantage of test time training to improve the geometric consistency of pointmaps. AerialMegaDepth [VGR25] instead focuses on aerial-ground geometric reconstruction from a data perspective.
@@
As a promising and powerful foundation for 3D reconstruction, VGGT [WCK25] presents a large feed-forward transformer-based architecture that directly predicts all essential 3D attributes, such as camera intrinsics and extrinsics, point maps, depth maps, and 3D point tracks, without the need for post-processing, leading to state-of-the-art 3D point and camera pose reconstruction. VGGT-Long [DTX25] extends VGGT to handle kilometer-scale sequences through a chunk-and-align pipeline, effectively mitigating its memory constraints in long-sequence handling.
@@
3D Gaussian Splatting (GS) [KKLD23] is a recent advance for efficient 3D reconstruction and rendering built on rasterization. Despite its high fidelity in reconstruction, 3DGS requires per-scene optimization, which limits its training efficiency and generalization capabilities. Recently, feed-forward 3DGS reconstruction methods have been developed by leveraging neural networks to directly predict Gaussian parameters, eliminating the need for per-scene optimization. We categorize these methods based on the representation of predicted Gaussian outputs: 2D map and 3D volume.
Figure 4: Representative frameworks with different outputs of 3D Gaussian representations, including Gaussian Map (e.g., Image Splatter and pointmap) and Gaussian Volume (e.g., Triplane and 3D Volume). The samples are adapted from [SRV24] [CXE24] [SZLP24] [ZYG24].